OpenStudy (anonymous):

Which of the following is equal to tan Ɵ + cot Ɵ ? (a) 1/ cot Ɵ + tan Ɵ (b) sin Ɵ + cos Ɵ/ (sin Ɵ) (cos Ɵ) (c) tan^2 Ɵ + 1 (d) (sec Ɵ) ( csc Ɵ) (e) sec^2Ɵ

OpenStudy (anonymous):

\[\frac{\sin(\theta)}{\cos(\theta)} + \frac{\cos(\theta)}{\sin(\theta)} \implies \frac{\sin^2(\theta) + \cos^2 (\theta)}{\sin(\theta) \cdot \cos(\theta)} \implies \frac{1}{\sin(\theta) \cdot \cos(\theta)} \implies cosec(\theta) \cdot \sec(\theta)\]

OpenStudy (anonymous):

what does it say after 1/ sin θ cosθ ?

OpenStudy (anonymous):

\[\frac{1}{\sin(\theta)} = cosec(\theta)\] \[\frac{1}{\cos(\theta)} = \sec(\theta)\]

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

Welcome dear..